The length of the rectangle is 1m less than twice it's width. What are the dimension of the rectangle. Perimeter is 40m
Answer by rfadrogane(214)

The length of the rectangle is 1m less than twice it's width. What are the dimension of the rectangle.Perimeter is 40m
Sol'n:
P = 2(L + W)---(1)
L = 2W - 1 ---(2)
substitute (2)into (1).
40 = 2(2W-1+W)
20 = 3W-1
21 = 3W
W = 7m
and for L:
L = 2(7)- 1
L = 13
thus, dimension are:
7m x 13m ---answer
^_^